OpenStudy (anonymous):

completing the square x^2 - 8x + 32 = 0

OpenStudy (anonymous):

(x-4)^2-32=0 (x-4)^2=32 x=4+-4*sqrt(2)

OpenStudy (anonymous):

\[x^2-8x=-32\] \[(x-4)^2=-32+16\] \[(x-4)=-16\] \[x-4=\pm\sqrt{-16}\] \[x-4=\pm4i\] \[x=4\pm4i\]

OpenStudy (anonymous):

typo in line 3 it should read \[(x-4)^2=-16\] sorry